two planes are 1680 miles apart and traveling toward each other. one place is traveling 80 miles per hour faster than the other plane. the planes meet in 1.75 hours. find the speed of each plane.
Thanks for helping me on this! :)

Answer by nerdybill(7384)

two planes are 1680 miles apart and traveling toward each other. one place is traveling 80 miles per hour faster than the other plane. the planes meet in 1.75 hours. find the speed of each plane.
.
Let x = speed of slower plane
then because "one plane is traveling 80 mph faster" we have
x+80 = speed of faster plane
.
From the "distance formula":
d = rt
where
d is distance
r is rate or speed
t is time
.
"distance traveled by slow plane" + "distance traveled by fast plane" = 1680
1.75x + 1.75(x+80) = 1680
1.75(x + (x+80)) = 1680
1.75(2x + 80) = 1680
1.75(2)(x + 40) = 1680
1.75(x + 40) = 840
(x + 40) = 840/1.75
x = 840/1.75 - 40
x = 480 - 40
x = 440 mph (speed of slower plane)
.
speed of faster plane:
x+80 = 440+80 = 520 mph